A solution is made by weighting 10g sodiumchloride and adding 40g of water this solution is:
a)10%(w/w)     b) 20%(w/w)   c) 25%(w/w) d) 5%(w/w)

Grade:12

1 Answers

Megha Sreejith
6 years ago
According to the equation of mass percentage , which is given in ncert, w/ w=(10 / 40 )(100)ie,w/w=25%